The average family size in the Philippines was reported as 4.25. A random of families in a particular street resulted in the following family sizes: 5, 6, 7, 4, 3, 2, and 9. At 𝛼=0.10, does the average family size more than the national average?


1
Expert's answer
2021-07-16T13:42:52-0400

Assume that the population is normally distributed.

"H_0:\mu=4.25"

"H_1:\mu>4 .25"

"\bar X =5.143"

"s=2.41"

"n=7"

"t=\frac{\bar X-\mu}{\frac{s}{\sqrt{n}}}"

"t=\frac{5.143-4.25}{\frac{2.41}{\sqrt{7}}}"

"=0.98"

"Cv=t_{0.1,6}=1.44"

Since the test statistic 0.98<1.44, we fail to reject the null hypothesis. The is no sufficient evidence to support the claim that the average family size is more than the national average of 4.25.